# Hongzhi Zang1, Mingjie Wei6,2, Si Xu3, Yongji Wu5, Zhen Guo3, Yuanqing Wang4,3, Hao Lin3, &Liangzhi Shi1, Yuqing Xie1, Zhexuan Xu1, Zhihao Liu7,2, Kang Chen4,2, Wenhao Tang1, &Quanlu Zhang3, Weinan Zhang6, Chao Yu1,2,†,∗, Yu Wang1,†

### v1

@@ Abstract
Recent progress in vision and language foundation models has significantly advanced multimodal understanding, reasoning, and generation, inspiring a surge of interest in extending such capabilities to embodied settings through vision-language-action (VLA) models. Yet, most VLA models are still trained with supervised fine-tuning (SFT), which struggles to generalize under distribution shifts due to error accumulation. Reinforcement learning (RL) offers a promising alternative by directly optimizing task performance through interaction, but existing attempts remain fragmented and lack a unified platform for fair and systematic comparison across model architectures and algorithmic designs. To address this gap, we introduce RLinf-VLA, a unified and efficient framework for scalable RL training of VLA models. The system adopts a highly flexible resource allocation design that addresses the challenge of integrating rendering, training, and inference in RL+VLA training. In particular, for GPU-parallelized simulators, RLinf-VLA implements a novel hybrid fine-grained pipeline allocation mode, achieving a 1.61x-1.88x speedup in training. Through a unified interface, RLinf-VLA seamlessly supports diverse VLA architectures (e.g., OpenVLA, OpenVLA-OFT), multiple RL algorithms (e.g., PPO, GRPO), and various simulators (e.g., ManiSkill, LIBERO). In simulation, a unified model achieves 98.11% across 130 LIBERO tasks and 97.66% across 25 ManiSkill tasks. Beyond empirical performance, our study distills a set of best practices for applying RL to VLA training and sheds light on emerging patterns in this integration. Furthermore, we present preliminary deployment on a real-world Franka robot, where RL-trained policies exhibit stronger generalization than those trained with SFT. We envision RLinf-VLA as a foundation to accelerate and standardize research on embodied intelligence.
Figure 1: Overview of RLinf-VLA. Through a unified interface, RLinf-VLA seamlessly supports diverse VLA architectures, multiple RL algorithms, and various simulators. To flexibly accommodate the integration of rendering, training, and inference in RL+VLA training, RLinf-VLA provides three GPU allocation modes: colocated, disaggregated, and a novel hybrid mode. A single unified model achieves 98.11% success on 130 LIBERO tasks and 97.66% on 25 ManiSkill tasks in simulation, while an RL-trained model demonstrates superior zero-shot generalization on a real-world Franka robot compared to the SFT model.
1 Introduction
@@
Vision-Language-Action (VLA) models represent a new generation of foundation models that aim to unify perception, language understanding, and embodied control (Ma et al., 2024; Firoozi et al., 2025). By leveraging vision-language models pretrained on internet-scale data and further training them on large, heterogeneous robot demonstration datasets (O’Neill et al., 2024; Khazatsky et al., 2024), VLAs can map raw sensor observations and natural language instructions directly into robot actions. This paradigm has already shown strong generalization across a wide range of tasks (Team et al., 2024; Kim et al., 2024; Liu et al., 2024; Wen et al., 2025; Shah et al., 2023; Black et al., 2024; Intelligence et al., 2025). However, the dominant training paradigm for VLAs remains supervised fine-tuning (SFT) via behavioral cloning on demonstration data (Kim et al., 2025). While effective in matching expert trajectories, SFT is highly vulnerable to distribution shift: small deviations from expert behavior can compound over time, pushing the policy into unfamiliar states and degrading robustness (Ross & Bagnell, 2010; De Haan et al., 2019; Belkhale et al., 2023; Foster et al., 2024). This mismatch between training and deployment distributions fundamentally limits scalability and generalization.
Reinforcement learning (RL) offers a complementary solution by directly optimizing cumulative task rewards through trial and error (Sutton & Barto, 1998). Unlike SFT, RL enables exploration beyond narrow expert data and equips policies with corrective strategies. Recent evidence suggests that RL fine-tuning can yield stronger out-of-distribution performance than SFT, particularly in both semantic alignment and execution robustness (Liu et al., 2025). Yet, progress on applying RL to VLAs remains small-scale (Liu et al., 2025) or fragmented (Lu et al., 2025; Li et al., 2025), primarily due to the lack of a unified and efficient framework to support large-scale experimentation. Unlike RL post-training for LLMs or VLMs, RL for VLAs requires repeated interactions with environments. These environments are typically realized through simulators, which also consume significant GPU memory and thus compete with model training and inference for memory resources—an issue that existing large-scale RL frameworks handle inefficiently. In addition, different works adopt diverse models, algorithms, and evaluation setups, making it difficult to compare methods fairly or distill general principles.
@@
To address these challenges, we present RLinf-VLA, a unified and efficient framework for scalable RL training of VLA models.
  • Unified: RLinf-VLA provides a cohesive interface that brings together multiple dimensions of RL for VLA training. It integrates different simulators, including ManiSkill (Tao et al., 2025) and LIBERO (Liu et al., 2023a), with features such as partial environment reset. It supports multiple algorithms, implementing both Proximal Policy Optimization (PPO) (Schulman et al., 2017; Ouyang et al., 2022) and Group Relative Policy Optimization (GRPO) (Shao et al., 2024; Guo et al., 2025). The framework is compatible with diverse VLA architectures, including OpenVLA (Kim et al., 2024) and OpenVLA-OFT (Kim et al., 2025) with action chunking. To flexibly accommodate the integration of rendering, training, and inference in RL+VLA training, RLinf-VLA provides three GPU allocation modes: colocated, disaggregated, and a novel hybrid mode. Switching between these modes can be achieved through straightforward configuration adjustments.
  • Efficient: Beyond unification, RLinf-VLA is designed for efficiency at both the system and algorithmic levels. It achieves a 2.27× improvement in throughput compared to the baseline framework. In particular, within GPU-parallelized simulators, the hybrid fine-grained pipelining mechanism further accelerates RL training, yielding a 1.61×–1.88× speedup. On the algorithmic side, RLinf-VLA incorporates several design choices, including lightweight critics, loss normalization, action masking, and rollout filtering, which collectively substantially enhance training efficiency.
Extensive experiments validate the effectiveness of RLinf-VLA. In simulation, a single unified model attains 98.11% success across 130 LIBERO tasks for the first time and 97.66% success across 25 ManiSkill tasks. More importantly, our study distills a set of best practices for RL-based VLA training:
  • PPO: action-level value estimation is superior to chunk-level estimation when using action chunks, and partial resets substantially improve sample efficiency.
  • GRPO: trajectory-length normalization and action masking are essential for stable training.
  • General Practices: larger rollout batch size is useful; LoRA may not directly affect performance but typically requires hyperparameter re-tuning.
Finally, we demonstrate preliminary deployment on a real-world Franka robot, where RL-trained policies exhibit stronger zero-shot generalization than those trained with SFT. Specifically, in a pick-and-place task with six previously unseen objects, the RL policy successfully completed 8 out of 30 trials, whereas the SFT policy failed in all 30 trials under the same zero-shot conditions. RLinf-VLA has been open-sourced and is maintained with frequent updates. We envision RLinf-VLA as a foundation to accelerate, standardize, and scale research in embodied intelligence.
Our contribution can be summarized as follows:
@@
  • Unified design: RLinf-VLA supports multiple simulators (ManiSkill, LIBERO), diverse VLA architectures (OpenVLA, OpenVLA-OFT), and RL algorithms (PPO, GRPO), with three execution modes, including a novel hybrid GPU allocation mode, enabling scalable and configurable training.
  • Efficiency in system and algorithm design: Hybrid fine-grained pipelining for GPU-parallelized simulators, colocated allocation for CPU-parallelized simulators, and algorithmic enhancements such as lightweight critics, loss normalization, action masking, and rollout filtering substantially improve training efficiency.
  • Strong empirical performance and generalization: A single unified model achieves 98.11% success on 130 LIBERO tasks and 97.66% on 25 ManiSkill tasks in simulation, while an RL-trained model demonstrates superior zero-shot generalization on a real-world Franka robot compared to the SFT model.
@@
  • Open-source and maintainable platform: RLinf-VLA is publicly released with frequent updates, providing a foundation to accelerate, standardize, and scale research in embodied intelligence.
2 Preliminary for Reinforcement Learning
@@
In this paper, we finetune VLA models with reinforcement learning to solve vision-based manipulation tasks, which we formulate as a partially observable Markov decision process (POMDP). A POMDP extends the standard MDP and is defined by the tuple $(\mathcal{S},\mathcal{A},P,r,\gamma,\Omega,O)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space (discrete or continuous, potentially chunked), $P(s^{\prime}\mid s,a)$ the state transition function, $r(s,a)$ the reward function, $\gamma\in[0,1]$ the discount factor, $\Omega$ the observation space, and $O(o\mid s)$ the observation function mapping a state $s$ to an observation $o$.
@@
We now recall several standard definitions in RL:
Definition 1 (Return).
$R_{t}=\sum_{k=0}^{\infty}\gamma^{k}r(s_{t+k},a_{t+k}).$ (2)
Definition 2 (Value Function).
The value function $V_{\pi}(s)$ is the expected return obtained from state $s$ at timestep $t$ when following policy $\pi$:
$V_{\pi}(s)=\mathbb{E}_{\pi}[R_{t}\mid s_{t}=s].$ (3)
Definition 3 (Action-Value Function).
The action-value function $Q_{\pi}(s,a)$ is the expected return obtained by executing action $a$ in state $s$ at timestep $t$ and thereafter following policy $\pi$:
$Q_{\pi}(s,a)=\mathbb{E}_{\pi}[R_{t}\mid s_{t}=s,a_{t}=a].$ (4)
Definition 4 (Advantage Function).
$A_{\pi}(s,a)=Q_{\pi}(s,a)-V_{\pi}(s).$ (5)
3 Framework Design
3.1 GPU Allocation Strategies
Figure 2: VLA+RL training loop
In robotic reinforcement learning, GPU resources are primarily consumed by the training process (Training) and the rollout process (Generation and Simulator). The training loop is shown in Figure 2. Simulators, however, vary by type: CPU-parallelized and GPU-parallelized simulators exhibit fundamentally different patterns of resource utilization. In CPU-parallelized simulators, the simulation logic is distributed across multiple CPU workers, while the GPU is mainly utilized for rendering and policy inference. As a result, CPU cores often become the primary bottleneck limiting overall performance. By contrast, GPU-parallelized simulators integrate rendering, physics simulation, and inference directly on the GPU, exploiting massive parallelism to achieve significantly higher throughput. However, this tight coupling also causes different components to compete for GPU memory and compute resources, making the overall performance highly sensitive to resource scheduling. Consequently, CPU-parallelized and GPU-parallelized simulators require distinct GPU allocation strategies, and efficient, flexible allocation across simulation, training, and inference is essential to maximize throughput.
(a) Colocated GPU Allocation
(b) Disaggregated GPU Allocation
(c) Hybrid GPU Allocation with Fine-Grained Pipelining
Figure 3: Different GPU allocation strategies.
To address this issue, our framework supports three allocation modes: colocated, disaggregated, and hybrid. Users can easily switch between these modes by simply modifying a YAML configuration file, greatly reducing the overhead of system customization. In particular, for GPU-parallelized simulators, we introduce a novel hybrid allocation with fine-grained pipelining strategy that further improves utilization efficiency. Figure 3 illustrates the core ideas of these allocation modes, while their algorithmic features are summarized here. For the full system design of RLinf, we refer the reader to Yu et al. (2025a).
3.1.1 Colocated Allocation
In this mode, all components (Training, Generation, and Simulator) share all GPUs simultaneously. This maximizes data parallelism and ensures that no GPU is left completely idle. Figure 3(a) provides an overview of the colocated allocation mode.
Offload Support
To make colocated allocation feasible under memory constraints, we introduce offloading support. Idle components can be temporarily moved to CPU memory, freeing GPU resources for the active ones. For Training and Generation, this offload–onload process is relatively straightforward. For the Simulator, however, we must manually store all essential environment states, so that the simulator can resume correctly once reloaded to the GPU.
Despite this support, colocated allocation is not always efficient. Two problematic cases often arise: 1) When multiple components reside on the same GPU, they may occupy memory concurrently even though only one is actively computing, leading to wasted capacity. 2) Although offloading reduces memory contention, its overhead can become substantial in practice. In particular, the frequent interaction between Generation and Simulator during rollout makes repeated offload–onload operations prohibitively expensive.
3.1.2 Disaggregated Allocation
@@
Figure 3(b) depicts the central idea of the disaggregated allocation mode. In this mode, each component is assigned to a distinct (possibly multi-GPU) partition, with no overlap across components. This ensures that every component can fully exploit its allocated resources. However, disaggregated allocation may still lead to GPU underutilization due to inter-component dependencies. For instance, the Simulator must wait for actions generated by the Generation, leaving some GPUs idle and creating “GPU bubbles.”
3.1.3 Hybrid Allocation with Fine-grained Pipelining
@@
To overcome the drawbacks of the above modes, we propose a hybrid strategy: hybrid allocation combined with fine-grained pipelining. In hybrid allocation, components can flexibly select GPUs. A typical configuration is to assign Generation and Simulator to different GPU partitions, while allowing Training to utilize all GPUs.
@@
On top of this, we introduce fine-grained pipelining to mitigate bubbles caused by inter-component dependencies. Specifically, a simulator instance on one GPU is partitioned into multiple sub-simulators, denoted as $S^{(1)},S^{(2)},\ldots,S^{(k)}$. The pipeline proceeds as follows [^1] In this section, superscripts denote simulator indices, while subscripts indicate timesteps.:

### v2

# Quanlu Zhang3, Weinan Zhang6, Chao Yu1,∗,†, Yu Wang1,†

@@ Abstract
Recent advances in vision-language-action (VLA) models have motivated the extension of their capabilities to embodied settings, where reinforcement learning (RL) offers a principled way to optimize task success through interaction. However, existing methods remain fragmented, lacking both a unified platform for fair comparison across architectures and algorithms and an efficient system design for scalable training. To address these challenges, we introduce RLinf-VLA, a unified and efficient framework for scalable RL training of VLA models. RLinf-VLA achieves unification by providing a unified interface that standardizes the integration of diverse VLA architectures, multiple RL algorithms, and heterogeneous simulators, enabling extensibility. To ensure efficiency, the system adopts a flexible resource allocation architecture for rendering, inference, and training workloads in RL pipelines. In particular, for GPU-parallelized simulators, RLinf-VLA introduces a hybrid fine-grained pipeline allocation strategy, yielding a 1.61x–1.88x training speedup. Using this unified system, models trained with RLinf-VLA demonstrate consistent performance improvements of approximately 20–85% across multiple simulation benchmarks, including LIBERO, ManiSkill, and RoboTwin. Furthermore, we distill a set of training practices for effective RL-based VLA training. We position RLinf-VLA as a foundational system to enable efficient, unified, and reproducible research in embodied intelligence.
I Introduction
@@
Vision-Language-Action (VLA) models represent a new generation of foundation models that aim to unify perception, language understanding, and embodied control [25, 9]. By leveraging vision-language models pretrained on internet-scale data and further training them on large, heterogeneous robot demonstration datasets [30, 16], VLAs can map raw sensor observations and natural language instructions directly to robot actions. This paradigm has demonstrated strong generalization across a wide range of tasks [40, 18, 23, 41, 34, 2, 15].
However, deploying VLA models typically requires post-training to mitigate distribution mismatch between training data and deployment environments. Without effective post-training, performance can degrade significantly when models encounter novel states, instructions, or execution conditions.
Reinforcement learning (RL) has emerged as an increasingly important post-training paradigm, as it directly optimizes cumulative task rewards through trial-and-error interaction [37]. In contrast to supervised fine-tuning (SFT), another commonly used post-training approach, RL enables exploration beyond narrow expert demonstrations and equips policies with corrective and adaptive behaviors. Recent studies indicate that RL fine-tuning can yield stronger out-of-distribution generalization than SFT, particularly in terms of semantic alignment and execution robustness [22].
Despite its promise, applying RL to VLA models remains largely small-scale or fragmented [22, 24], making the development of new RL algorithms for VLAs expensive and cumbersome. Although SimpleVLA-RL [20] enables large-scale RL training for VLAs by building on VeRL [36], an RL codebase originally designed for large language models, it lacks system-level optimizations tailored to embodied settings, where simulators compete with model inference and learning for GPU resources. This limitation constrains scalability and substantially increases training time.
Moreover, online RL requires repeated and tightly coupled model–environment interactions. Without a system design explicitly tailored to this execution pattern, significant GPU idle time and pipeline bubbles can arise, leading to inefficient resource utilization. Finally, existing works often rely on disparate model architectures, RL algorithms, and evaluation protocols, making fair comparison difficult and hindering the extraction of general principles for RL-based VLA training.
@@
To address these challenges, we present RLinf-VLA, a unified and efficient framework for scalable reinforcement learning of vision-language-action models. Our main contributions are summarized as follows:
@@
  • Unified system abstraction. RLinf-VLA provides a unified interface that supports multiple robotic simulators (ManiSkill [39], LIBERO [21], RoboTwin [8]), diverse VLA architectures (OpenVLA [18], OpenVLA-OFT [17]), and reinforcement learning algorithms (PPO [33], GRPO [35]). The system exposes three execution modes, including a novel hybrid GPU allocation mode, enabling scalable and configurable training across heterogeneous setups.
  • Efficient system and algorithm design. RLinf-VLA introduces hybrid fine-grained pipelining for GPU-parallelized simulators and collocated execution for CPU-parallelized simulators, substantially improving training throughput, with speedups of up to 2.27×. In addition, we incorporate a set of algorithmic optimizations that further enhance training efficiency and stability.
  • Strong empirical performance and generalization. Using a single unified model, RLinf-VLA achieves a 98.11% success rate on 130 LIBERO tasks and 97.66% on 25 ManiSkill tasks. RLinf-VLA also achieves an average success rate of 84.63% on six RoboTwin tasks, realizing an average performance improvement of 63.75% on these tasks, which demonstrates its powerful performance and generalization capabilities in post-training.
@@
  • Open and extensible platform. RLinf-VLA is released as an open-source and actively maintained platform, providing a practical foundation to accelerate, standardize, and scale reinforcement learning research for embodied intelligence.
II Related Work
II-A Reinforcement Learning for VLA
Vision-Language-Action models (VLAs) [17, 18, 2, 4, 15, 43, 5, 1, 46, 6] represent a multimodal foundation model architecture designed for robotics. RL is increasingly adopted to address the limitations of static imitation learning in VLAs. GRAPE [44] employs Direct Preference Optimization (DPO [31]) on partial trajectories to align models with human intent. Iterative approaches [12, 38] combine SFT with RL stages to balance stability and performance. VLA-RL [24] utilizes Proximal Policy Optimization (PPO [33]) to exhibit significantly stronger generalization to unseen objects and environments compared to their SFT counterparts [22]. SimpleVLA-RL [20] and related variants [38] apply Group Relative Policy Optimization (GRPO [35]) to improve training efficiency and performance without complex reward design. $\pi_{RL}$ [7] implements the policy gradient algorithm to flow-matching model for the first time. $\pi^{*}_{0.6}$ [14] introduces RECAP to train a binarized value function and VLA. A unified and efficient infrastructure is essential for advancing VLA via RL, yet current implementations remain fragmented [22, 24]. Although frameworks such as SimpleVLA-RL [20] have emerged, they largely follow generic LLM RL post-training paradigms and fail to adequately account for the system-level optimizations and interface extensibility required by embodied intelligence tasks.
II-B Simulators
Simulators [19, 26, 27, 29, 28, 10, 3] offer scalable, safe, and controllable environments that bypass physical hardware constraints. They facilitate massive parallel data generation and automated resets, essential for training and evaluating generalizable VLA policies. Notable platforms include ManiSkill [39] for physics-rich manipulation, LIBERO [21] for instruction-conditioned reasoning, and RoboTwin [8] for bimanual tasks with domain randomization. Consequently, the construction and utilization of simulators are indispensable for RL. However, inconsistent interfaces across simulators often necessitate extensive code restructuring. A unified interface is therefore critical to enable seamless switching between diverse simulation benchmarks.
III Preliminary
III-A Reinforcement Learning
@@
To guide the design of our efficient training infrastructure, we first formalize the problem of finetuning VLA models using Reinforcement Learning (RL). We model vision-based manipulation tasks as a partially observable Markov decision process (POMDP). A POMDP extends the standard MDP and is defined by the tuple $(\mathcal{S},\mathcal{A},P,r,\gamma,\Omega,O)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space (discrete or continuous, potentially chunked), $P(s^{\prime}\mid s,a)$ the state transition function, $r(s,a)$ the reward function, $\gamma\in[0,1]$ the discount factor, $\Omega$ the observation space, and $O(o\mid s)$ the observation function mapping a state $s$ to an observation $o$.
@@
III-B Pipeline of Reinforcement Learning for VLA models
Figure 1: Pipeline of Reinforcement Learning for VLA models
The RL pipeline for VLA models comprises three distinct components: Generation, Simulator, and Training, as illustrated in Figure 1. GPU resources are primarily allocated between the rollout phase (Generation and Simulator) and the optimization phase (Training). A key distinction from LLM-based RL lies in the hierarchical action formulation. While LLMs typically operate on single tokens [35], VLA policies often predict a single action [18] or a chunk of actions to ensure smooth control [45]. In our framework, we unify these representations by treating a single action as a chunk of size one. Crucially, this introduces a three-level hierarchy: Chunk $\to$ Atomic Action $\to$ Token. Each atomic action comprises multiple tokens, where each token typically corresponds to a specific dimension of the robot’s action space (e.g., end-effector pose or joint angles). Moreover, the integration of simulators imposes substantial computational overhead on the generation process, demanding significant CPU and GPU resources for physics simulation and rendering. The interaction typically proceeds as follows [17]: the Generation module infers an action chunk based on the current observation; the Simulator executes this chunk and returns the observation from the final timestep. This loop continues until trajectory collection is complete, after which the Training module updates the policy using the gathered data.
IV Method
Section IV-A introduces flexible GPU Allocation Strategies for improved resource scheduling across different simulator types. Section IV-B presents a Unified Interface that seamlessly integrates diverse simulators, VLA models, and RL algorithms. Building on this system design, Section IV-C distills key Design Choices for scalable RL training of VLA models.
IV-A GPU Allocation Strategies
The RL training pipeline (described in Section III-B) for VLA involves varying resource demands from Training, Generation, and Simulator. Resource bottlenecks depend heavily on the simulator type: CPU-parallelized simulators are typically CPU-bound, using GPUs primarily for rendering and inference, whereas GPU-parallelized simulators execute simulation, rendering, and inference entirely on the GPU. While the latter offers higher throughput, it creates severe contention for GPU memory and compute. To maximize efficient utilization across these diverse setups, flexible and optimized GPU allocation strategies are essential. Our framework supports flexible and easily configurable allocation modes: collocated, disaggregated, and a novel hybrid mode.
IV-A1 Collocated Allocation
Figure 2: Collocated Mode.
In this mode (Figure 2), all components co-exist on the same set of GPUs. The original version of collocated mode aimed to maximize GPU utilization by maintaining only one component on all GPUs at any given time. When a component finished its computation, it would be offloaded from GPU to CPU memory, and then onloaded again when needed for its next task. However, under the embodiment setting, both Simulator and Generation require GPU resources and interact frequently in an iterative manner. The overhead introduced by repeated offload-onload operations at each interaction becomes prohibitively large and unacceptable.
Therefore, we implement a modified collocated strategy where offload and onload operations for the Simulator and Generation occur only at the beginning and end of the rollout phase, as shown in the figure. While this approach successfully avoids the frequent offload-onload overhead during rollout, it cannot fully utilize GPU resources and remains sub-optimal. Since Generation and Simulator must interact iteratively, they spend significant time waiting for each other, occupying GPU memory without performing computation, leading to resource wastage and limited scalability.
IV-A2 Disaggregated Allocation
Figure 3: Disaggregated Mode (see Figure 2 for legend).
@@ IV-A2 Disaggregated Allocation
In this mode, each component is assigned to a distinct (possibly multi-GPU) partition, with no overlap across components. Figure 3 depicts the central idea. This ensures that every component can fully exploit its allocated resources. This mode is simple to implement, but it lead to GPU underutilization due to inter-component dependencies. For example, in Figure 3, the gpu 4, 5 are completely idle during the rollout phase util training.
IV-A3 Hybrid Allocation with Fine-grained Pipelining
Figure 4: Hybrid Mode (see Figure 2 for legend).
@@ IV-A3 Hybrid Allocation with Fine-grained Pipelining
To overcome the drawbacks of the above modes, we propose a hybrid strategy: hybrid allocation combined with fine-grained pipelining. In hybrid allocation, components can flexibly select GPUs. A typical configuration is to assign Generation and Simulator to different GPU partitions, while allowing Training to utilize all GPUs. Figure 4 shows an illustration. However, the resources remain underutilized. For instance, the Simulator must wait for actions generated by the Generation, leaving some GPUs idle and creating “GPU bubbles”.
Figure 5: Hybrid Mode with Fine-grained Pipelining (see Figure 2 for legend).
@@
On top of this, we introduce fine-grained pipelining to mitigate bubbles caused by inter-component dependencies. Specifically, a simulator instance on one GPU is partitioned into multiple sub-simulators, denoted as $S^{(1)},S^{(2)},\ldots,S^{(k)}$. The pipeline proceeds as follows, where superscripts denote simulator indices and subscripts indicate timesteps: